200 TICKETS SOLD, ADULTS = X = $8, CHILDREN = 200-X = 5$, TOTAL = 1441, HOW MANY CHILDREN?
Answer by mananth(16946)

200 TICKETS SOLD, ADULTS = X = $8, CHILDREN = 200-X = 5$, TOTAL = 1441, HOW MANY CHILDREN?
Adult ticksts cost $8
Number of adults = X
Priice of adult tickets = 8x
.
Children tickets cost $5
Number of children = 200-x
Price for children tickets = 5(200-x)
8x+5(200-x)= 1441
8x + 1000-5x= 1441
3x=1441-1000
3x=441
x= 147 number of adults
200 -47 = 153 number of children